An amount of R3 000,00 is invested for three years at a simple interest rate and it earned R905,00
interest. Determine the simple yearly interest rate at which the money was invested. Give your answer
as a percentage, rounded to two decimal digits.
1
Expert's answer
2021-06-15T05:35:11-0400

P=R3000.00

T=3 years

SI=R905


"SI=P\times R\times T\\R905=R3000\times \frac{r}{100}\times 3"


"R905=\frac{R9000r}{100}"


"R905=\frac{R905}{R90}"


"r=\frac{R905}{R90.}"


"r=10.06\%"